2 Find (4725, 3234) and (4725, 3234).
Answer #1

2. Find (4725,3234) $ [4125,3234] 2010-0 (6725, 8232) Here a=4725, b=3234 4725= 3234 (1) + 1491 3234 = 1491(2) +252 1491 = 25